i don't really know which category this goes under but here goes nothing...
an object that is projected straight downward with initial velocity v feet per second travels a distance s=vt+16t^2, where t=time in seconds. If Ramon is standing on a balcony 84 ft above the ground and throws a penny straight down with an initial velocity of 10 feet per second, in how many seconds will it reach the ground
a. 2 seconds
b 3 seconds
c 6 seconds
d 8 seconds

Answer by ankor@dixie-net.com(22740)

an object that is projected straight downward with initial velocity v feet per second travels a distance s=vt+16t^2, where t=time in seconds.
If Ramon is standing on a balcony 84 ft above the ground and throws a penny straight down with an initial velocity of 10 feet per second, in how many seconds will it reach the ground
:
s = 84 ft, v = 10 ft/sec
Rewrite the equation to
16t^2 + 10t = 84
16t^2 + 10t - 84 = 0
simplify divide by 2
8t^2 + 5t = 42
You can use the quadratic formula to find t, but this will factor to
(8t+21)(t-2) = 0
the positive solution is all we want here
t = 2 seconds
:
:
Check this in the original equation
s = vt + 16t^2
v=10, t=2
s = 10(2) + 16(2^2)
s = 20 + 64
s = 84 ft